OpenStudy (anonymous):

factor x^2-15+54 pleaase answer asap

OpenStudy (anonymous):

Did u mean x^2-15x+54?

OpenStudy (anonymous):

I assume that middle term is -15x. You're looking for two numbers such that: a + b = -15 a * b = 54 Those two numbers will be in your factored form as: (x+a)(x+b)

OpenStudy (anonymous):

yes roshan

OpenStudy (anonymous):

69 is hot

OpenStudy (anonymous):

If it is then, x^2-9x-6x+54 x(x-9)-6(x-9) (x-9)(x-6)

OpenStudy (anonymous):

grrrr

OpenStudy (anonymous):

thank u

OpenStudy (anonymous):

(x-6)(x-9) = grrrrrr

OpenStudy (anonymous):

I agree.

OpenStudy (anonymous):

imani, if you just want answers, please use wolframalpha.com and don't take up tutors time/resources.